Tech Staff Meeting Notes February 21, 2006
Current Top Priorities are:Station Maintenance, Broadband Upgrade, NC Node
There are four levels of priority.
|Priorities||Wish to try to get these done (or make progress on) before the next meeting (yeah right)|
|Second Tier||Fit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month|
|As Time Permits||Fit these in when you can but not instead of first or second priorities|
|Other||Stuff we hope to get to and don't want to forget about|
This redistributed task list agreed to at the November meeting. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.
Greg's February ToDo
is incorporated into this rest of this page.
BB network upgrade.
Steve has the software running to generate the McNamara PDF plots. Still need the dataless. Strict determination of physical ground motion units is not required for this test. We'll use a generic dataless seed.
Need to track timing.
Need to develop a telemetry plan. Ideally stations won't move but they can if they have to. Ideally they'll go to a node but can go to local internet POP if necessary.
James and Bob weren't able to be at the meeting. These are notes from January.
Columbia, TN permission is a wash. Still don't have an eastern hard rock station.
Need to notify James when we do the fiber isolation at the CVTN strongmo station so he can verify we didn't muck up the GPS station.
Camp Robinson needs a routine service. Contact info needs to be updated and is likely best done on site.
Need to test aerocom telemetry path at Steele. Station is at private business near airport and the plan is to telem to High School on the other side of town where we can get internet.
Piggot not yet converted to ethernet though permission is secured. IT tech is difficulty to contact.
Need a new plan for MS County Airport at East Prairie, MO.
Reelfoot needs to be converted from Solar to AC.
Reviewed plan. The map shows stations
with yellow triangles being new ones.
Agreement with Blue Ridge is complete. John will bring the LRNC
node to Memphis the week of April 17 for refurbish and redeployment
to Mt. Mitchell.
HAIL was swapped with a 2g, need to swap SEAR
Continued problems with mss100's losing connectivity (SMOC: simple matter of connectivity).
SWET-3H52, ADSC-5657, and CSTN-5637 need to have the leap second corrected. SHTN-5648 and TRSC-5662 might also but need a site visit to verify.
Need to upgrade trimble firmware to avoid leap second error next time. A workable upgrade procedure doesn't yet exist. Jim, Wurth, and DTA or working on it.
CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. James brought back some nice pictures from CVTN to help figure out how to do it (e.g. the conduit is way too small).
Routine Maintenance. Jim knows what to do.
Emergency lights in grad building don't work except for the ones in the hallway.
Second Tier Priorities
Gordana Vlahovic is getting a cmg6td to install at North Carolina
Central University. John will pay a visit to provide guidance on
Bran Potter sent a list of power options. Alena gave some specs for
solar operation. Looks like the best way is to do a low voltage feed
like TZTN. Should we use VSAT? Still need to somehow provide real-time
wiggle display in the Geology building.
Steiner Brain Dump Seminar
Beginning this summer, Greg will be giving a weekly non-credit seminar.
It will primarily be for the tech staff and will cover all the knowledge
we need to pass from Greg to other staff. Others at CERI will be welcome
to participate if they wish.
New EW and Hydra responsibilities require more space for testing and
configuring computer hardware. This new space will be in 3904/211.
The copier will go where the light table is and the light table
will go where ever JM Chiu wants it to.
Brent Rosenblad at UofMO, Columbia will be doing surface wave analysis
with a large vibroseis source in the NMSZ. Jim is coordinating to see
if tests near a station are possible.
FDL. Bill is doing a great job of getting it organized. We have all three Dell latitude laptops accounted for. One is ready to go in Horton's office, one is being used by Bodin, and the third was cannabilized for parts. Mitch and Paul are working on getting replacements.
As Time Permits
This project is in O&M mode--no more development. Al will need to clean off the recorder disks every other week. We can't afford to spend anymore time improving at the expense of other projects. Hence we will attempt to keep it going as best we can. Doesn't work too bad given all the possible points of failure.
Got the modem and internet works. Need to integrate into CERI central processing.
MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data.
VBMS. Vault installation went out for bid.
Langston student projects continue. See Greg's list above.